Broward Manufacturing recently reported the following information:

Net income $515,000
ROA 12%
Interest expense $195,700
Accounts payable and accruals $1,050,000

Broward's tax rate is 40%. Broward finances with only debt and common equity, so it has no preferred stock. 40% of its total invested capital is debt, while 60% of its total invested capital is common equity. Calculate its basic earning power (BEP), its return on equity (ROE), and its return on invested capital (ROIC). Round your answers to two decimal places.

BEP %
ROE %
ROIC %
Answer #1

BEP = EBIT/ Total assets

EBIT = Net income+ Taxes + Interest

= Net income + Net income*40/60 + 195700

= 515000+ 515000*40/60+ 195600

=1053933.33

Total assets =Net Income/ ROA

= 515000/12%

= 4291667

BEP = 1053933.33/ 4291667

= 24.56%

2: Equity = 60%*Total assets = 60%*4291667= 2575000

ROE = Net income/ Equity

= 515000/2575000

= 20%

3: ROIC = NOPAT/ Invested capital

=(EBIT*(1-tax))/ (Total assets- accounts payable and accruals)

= (1053933.33*0.6)/( 4291667- 1050000)

=19.51%